OpenStudy (anonymous):

James wants to work up to doing 40 sit-ups a day. He plans to do 5 sit-ups the first day, 9 sit-ups the second day, 13 sit-ups the third day, and so on. On what day will he do 45 sit-ups?

OpenStudy (cwrw238):

45 = 5 + (n - 1) 4 solve for n will give you the day

OpenStudy (cwrw238):

day 1 5 2 9 3 13 4 17 and 45-17 = 28 more = 7 days total 11 days

OpenStudy (cwrw238):

ok ? this is an arithmetic series with common difference 4 nth term = a + (n-1)d where a = first term and d = common difference thats where the equation 45 = comes from